User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Meeting 30th March

Chair: Bryn

Present: Jim, Lucia, Liz, Tiege, Jamie, Jonny, Ross, Arwin, Lucas, Ellen

Apologies: Jesse

Absent: Nick

General Issues:

  • Revisiting 5pm-finish shifts
    • Tell Ellen THIS WEEK if you are okay with the 5pm till finish shifts.
    • By default we aren't scheduling 5 till finish anymore - swap if you want to do them.
  • How are people with their shift loading?
    • Update availability if you want fewer shifts
  • Bed is starting to gross me out
    • If you use it, please take it home and wash it.
  • Free your calendars for Nov 28 - Dec 12. Continuous observing.
  • If the power is out on the plugs in the observation room - check the switchboard for any tripped switches.

Handover Issues:

  • RD1704 VC rack timing out - had to travel out to Mt. P to reset the VC rack and the TTY Distributor.
    • Due to power failure
    • Strongly recommended to do setup on the previous day.
  • equip.ctl file check?
  • Wind sensor at hb locked up - continuous wind stow. Do we have a fix for this in the wiki?
    • Ke & Yg (?) have IP switches - Requires turning off and on again.
    • Anemometer locked up - restart (standard monica restart process)
  • Network issues at Ke - R1784
    • Lost communication to VPN and Ke network. CMOS battery died in pcfske, rebooted, but wouldn't reconnect to network. Had to reboot pcfske through mk5ke. Might need to follow up on network configuration for the pcfske machine.
    • If you can't ping a machine from outside a site, try pinging from something else on the local network.
  • Sending off logs - Ross had problems sending off the logs for R1785. Flogit backup script can be used if you have issues with the ftp to France.
/home/www/auscope/opswiki/data/pages/operations/agenda2017_03_30.txt · Last modified: 2017/03/30 02:35 by Bryn Emptage